HUBBLE SPACE TELESCOPE DAILY REPORT #3054 PERIOD COVERED: 0000Z (UTC) 02/11/02 - 0000Z (UTC) 02/12/02 Daily Status Report as of 043/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of STIS/CCD 9088 (Next Generation Spectral Library of Stars) The Space Telescope Imaging Spectrograph (CCD) was used to produce a "Next Generation'' Spectral Library of 600 stars for use in modeling the integrated light of galaxies and clusters by using the low dispersion UV and optical gratings of STIS. The library will be roughly equally divided among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling the entire HR-diagram in each bin. Such a library will surpass all extant compilations and have lasting archival value, well into the Next Generation Space Telescope era. As described in yesterday's report and HSTAR 8496, the acquisition for the first iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting one observation. Then, per HSTAR 8497 and 2.1, the acquisition for the third iteration of this proposal failed, the take data flag remained down and all observations were lost. No other problems occurred. 1.2 Completed Eleven Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. As described in yesterday's report and HSTAR 8496, the acquisition for the first iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting one observation. Then, per HSTAR 8497 and 2.1, the acquisition for the final iteration of this proposal failed, the take data flag remained down and all observations were lost. There were no other reported problems. 1.3 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.4 Completed STIS/CCD 9180 (Gamma-ray Burst Progenitors: Probing Their Environment) The Space Telescope Imaging Spectrograph (CCD) was used to perform a target of opportunity observation of gamma ray burster (GRB), GRB-020124. GRB astronomy is a field maturing at a phenomenal rate. Three important new observational and theoretical discoveries, formulated over the last twelve months, allow the proposer to address new, and in many cases, more sophisticated questions than could have been posed previously. These developments: the discovery of X-ray lines in GRB 991216; the observation that N_H as deduced from X-ray afterglow are one to two orders of magnitude larger than the dust extinction inferred from optical afterglow; and the growing realization that the afterglow emission may exhibit features of dust echoes, appear to offer unexpected and new diagnostics that will directly inform us about the progenitor, the circum-progenitor material and the immediate interstellar environs. There were no reported problems. 1.5 Completed WF/PC-2 8941 (Cycle 10 UV Earthflats) The WF/PC-2 was used to monitor flat field stability by obtaining sequences of earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set. The proposal had no problems. 1.6 Completed Two Sets of WF/PC-2 8940 (Cycle 10 Earth Flats) The WF/PC-2 was used to monitor flatfield stability by obtaining sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The proposal completed without incident. 1.7 Completed Three Sets of WF/PC-2 8942 (Cycle 10 Intflat Sweeps and Linearity Test) The WF/PC-2 was used to monitor the pixel-to-pixel flatfield response and provide a linearity check. These intflat sequences are done once during the year and the images will provide a backup database in the event of complete failure of the visflat lamp as well as allow monitoring of the gain ratios. There were no reported problems. 1.8 Completed Five Sets of WF/PC-2 8935 (Cycle 10 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported problems. 1.9 Completed WF/PC-2 8939 (Cycle 10 Internal Monitor) The WF/PC-2 was used to calibrate the internal monitor, to be run weekly to monitor the health of the cameras. No problems were encountered. 1.10 Completed STIS/CCD/MA1/MA2 9161 (The Ultraviolet Properties of Obscured QSOs) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to observe a subset of The Two Micron All-Sky Survey {2MASS} -discovered QSOs to: 1} determine if UV absorption- line features are present as seen in several other highly polarized AGN populations; 2} test relationships between X-ray and UV/optical attenuation, AGN spectral type, optical polarization, and detected UV absorption-line features; and 3} characterize the spectral energy distributions and UV properties of obscured QSOs. There were no problems. 1.11 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 1.12 Completed STIS/CCD/MA1 9067 (UV Detectability of Bright Quasars in the Sloan Fields) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to take MAMA spectra of approximately 30 new, bright, high-redshift quasars in each of the next three cycles. The observations completed with no reported problems. 1.13 Completed STIS/CCD 9114 (SINS: The Supernova Intensive Study-- Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to perform closer observations of supernovae that will create the chemical history of the Universe, energize the interstellar gas, stiffen the spine of the extragalactic distance scale and provide the only evidence for an accelerating universe. A violent encounter is underway between the fast-moving debris and the stationary inner ring. Monitoring this interaction will help solve the riddles of stellar evolution posed by the enigmatic three-ring system of SN 1987A. Our UV observations of Ly- Alpha emission reveal the present location and velocity of a remarkable reverse shock that provides a unique laboratory for studying fast shocks and a powerful tool for dissecting the structure of the vanished star. No problems were encountered. 1.14 Completed STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 6 Per HSTAR 8497, the acquisition scheduled for 042/234408Z failed when the scan step limit was exceeded on FGS-3. The proposals described in 1.1 and 1.2 were affected. Scheduled Re-acquisitions: 9 Successful: 9 2.2 FHST Updates: Scheduled: 15 Successful: 15 2.3 Operations Notes: Joint Integrated Simulation (JIS) #6 (Rendezvous) executed yesterday from 9:30 a.m. until 8:00 p.m. The 486 engineering status buffer (ESB) was dumped and cleared at 042/1148Z per ROP DF-18A. Using the same ROP, the ESB limit for SESBSLD (=3) was set at 043/0001Z The ephemeris table was uplinked at 043/0149Z as directed by ROP DF-07A. Due to the acquisition failure described in 2.1, the STIS flight software error counter was cleared at 043/0150Z (ROP NS-12). Using ROP SR-1A, the SSR-1 EDAC error counter was cleared at 043/0224Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: JIS #5 (EVA-4) is scheduled to be conducted today from 9:30 a.m. until approximately 6:30 p.m. Continuation of normal science observations and calibrations. /CAW